# SlideChat: A Large Vision-Language Assistant for Whole-Slide Pathology Image Understanding

### v2

@@ A.1 Data Source
A.2 Curation Scope and Prompt
A.2.1 Scope
Microscopy
Diagnosis
Clinical
A.2.2 Designed Prompts
Report Cleaning Prompt.
Table 6: Prompts for report clean and caption generation.
[Report Clean Prompt]
[Caption Generation Prompt]
[System Prompt]
[General Prompt]
[Objective Prompt]
Table 8: Prompt for Converting Labels into QA Pairs
[Label Transformation Prompt]
Report Cleaning Prompt.
The prompt used to clean up the report from the original TCGA report is represented in Table 6. This process effectively eliminates extraneous noise from the report, thereby establishing a more solid foundation for caption and QA pairs generation.
Caption Generation Prompt.
The prompt used for caption generation from the refined report is detailed in Table 6, ensuring that the generated caption effectively captures essential summarized information in report.
Question-Answers Generation Prompt.
The prompt used to extract QA from reports mainly consist of 4 parts (i.e., <Cleaned Report>+ System Prompt + Objective Prompt + General Prompt), and the detailed content of each part is illustrated in Table 7
Label Transformation Prompt.
The prompt for transforming BCNB dataset is illustrated in Table 8. We employ GPT to transform individual labels into a question-answer format based on the task type and corresponding classification labels, facilitating the testing of MLLM. For instance, in the context of a tumor type classification task, <Task>represents “Tumor Type”, while <label 1>, <label 2>, and <label 3>are “Invasive ductal carcinoma”, “Invasive lobular carcinoma”, and “Other Type”, respectively, enabling the generation of relevant QA pairs.
A.3 Data Statistics
We have compiled statistics on the number of VQA instances for each category within SlideBench VQA (TCGA) in Table 9. Each subcategory contains over 500 VQA instances, ensuring a robust representation across all areas, which supports comprehensive model evaluation and facilitates in-depth performance analysis. We provide an overview of the sample sizes and detailed original label information for the seven classification tasks within the BCNB dataset in Table 10.
@@ A.3 Data Statistics
A.4 Multimodal Dataset Comparsion
Recently, several multimodal pathology datasets have been introduced for pathology applications. However, these datasets are often constrained in both scope and scale, as they primarily focus on either patch-level analysis or limited available data. In contrast, our proposed SlideInstruction and SlideBench, provided as open-source resources, significantly expand the dataset size while enhancing its versatility, as shown in Table 11.
@@ A.4 Multimodal Dataset Comparsion
Dataset Level Data Type Curation Type Scope Number # Availability
PathChat (Lu et al., 2024b) Patch Patch and Q/A pairs Human+GPT - 257,004
Quilt-Instruct (Seyfioglu et al., 2024) Patch Patch and Q/A pairs GPT - 107,131
WSI-VQA (Chen et al., 2024a) Slide WSI and Q/A pairs GPT - 8,672
PathText (Chen et al., 2023) Slide WSI-Caption pairs GPT - 9,009
HistGen (Guo et al., 2024) Slide WSI-Reports pairs GPT - 7,753
Prov-Path (Xu et al., 2024a) Slide WSI-Reports pairs GPT - 17,383
CR-PathNarratives (Zhang et al., 2023) Slide WSIs with multi-modal annotations Human - 174
PathAlign (Ahmed et al., 2024) Slide WSI-Reports pairs Human - 354,089
@@
Appendix B Experiment
Figure 9: Comparing model outputs on SlideBench.
B.1 Implementation Details
We preprocessed each WSI by segmenting it into 224 × 224 nonoverlapping patches at a 20× magnification level, excluding background regions. We implemented our model using the Xtuner (Contributors, 2023) toolkit and trained it across two stages on 8 × NVIDIA A100 GPUs. The training process consists of an alignment phase followed by instruction fine-tuning: Stage 1: We freeze the LLM and train the Projection and Slide Encoder with WSI-caption data for 3 epochs, using a learning rate of 0.001. Stage 2: We unfreeze the LLM, Slide Encoder, and Projection, training the model on WSI instruction-following data for 1 epoch, with a learning rate of 0.00002. Both stages are optimized using AdamW.
B.2 Ability Showcase
B.2.1 Captioning Ability
The examples shown in Figure 7 illustrate the capability of our model, SlideChat, to effectively perform whole-slide image captioning tasks. SlideChat demonstrates its proficiency in generating detailed and contextually accurate summaries for complex pathological whole-slide images, accurately capturing key clinical findings and pathological features. Whether summarizing broad findings, explaining pivotal details, or highlighting core results, SlideChat showcases an advanced understanding of whole-slide images, providing concise yet informative reports that align with clinical terminology and expectations.
B.2.2 VQA Ability
Figure 8 showcases the conversational examples of SlideChat, demonstrating its ability to accurately answer a range of questions based on WSIs, covering diverse aspects such as histological classifications, tumor grading, lymph node involvement, and treatment decisions. SlideChat effectively interprets complex pathological data, engages in nuanced question-and-answer exchanges, and delivers clinically relevant responses. This reflects its potential as an intelligent assistant capable of supporting pathologists in diagnostic decision-making by providing insightful, context-aware dialogue grounded in visual pathology data.
B.2.3 Comparing Model Outputs
Figure 9 presents a comparative analysis of the outputs from SlideChat and other models within SlideBench. The examples illustrate SlideChat’s remarkable capacity to precisely classify tumors, identify distinct histological features, and describe the structural organization of tumor cells from WSIs. SlideChat demonstrates a unique proficiency in capturing both local and global features—seamlessly integrating detailed microscopic characteristics with broader contextual understanding to deliver accurate and clinically meaningful interpretations. In contrast, existing models are limited to processing small pathology images, often yielding ambiguous or incorrect classifications. This underscores SlideChat’s advanced capability in comprehending whole-slide images by incorporating both intricate details and a comprehensive visual perspective.
B.3 Detailed Test Performance
@@
LLaVA-Med Patch 50.04 40.63 40.38 56.95 47.34
GPT-4o 37.07 38.76 39.93 37.60 38.28
@@
LLaVA-Med Patch 37.25 28.57 30.41 20.71 47.27 32.78
GPT-4o 22.95 26.76 18.06 21.06 27.82 23.10
@@
LLaVA-Med Patch 62.04 53.98 53.04 26.54 47.96
GPT-4o 50.00 50.08 44.16 32.64 43.42
@@
LLaVA-Med Patch 23.95 36.62 40.19 50.76 23.72 18.99 15.05 30.10
GPT-4o 0 0 0 0 0 0 0 0
@@
B.3.1 Performance on SlideBench-VQA (TCGA)
@@
B.3.2 Performance on SlideBench-VQA (BCNB)

### v3

@@ 6.1 Data Source
6.2 Data Statistics
We have compiled statistics on the number of VQA instances for each category within SlideBench VQA (TCGA) in  Tab. 5. Each subcategory contains over 500 VQA instances, ensuring a robust representation across all areas, which supports comprehensive model evaluation and facilitates in-depth performance analysis. We provide an overview of the sample sizes and detailed original label information for the seven classification tasks within the BCNB dataset in  Tab. 6.
@@ 6.2 Data Statistics
6.3 Curation Scope and Prompt
6.3.1 Scope
Microscopy
Diagnosis
Clinical
6.3.2 Designed Prompts
Report Cleaning Prompt.
Report Cleaning Prompt.
The prompt used to clean up the report from the original TCGA report is represented in Tab. 7. This process effectively eliminates extraneous noise from the report, thereby establishing a more solid foundation for caption and QA pairs generation.
Table 7: Prompts for report clean and caption generation.
[Report Clean Prompt]
[Caption Generation Prompt]
Caption Generation Prompt.
The prompt used for caption generation from the refined report is detailed in Tab. 7, ensuring that the generated caption effectively captures essential summarized information in report.
Question-Answers Generation Prompt.
The prompt used to extract QA from reports mainly consist of 4 parts (i.e., <Cleaned Report>+ System Prompt + Objective Prompt + General Prompt), and the detailed content of each part is illustrated in Tab. 8
[System Prompt]
[General Prompt]
[Objective Prompt]
Label Transformation Prompt.
The prompt for transforming BCNB dataset is illustrated in Tab. 9. We employ GPT to transform individual labels into a question-answer format based on the task type and corresponding classification labels, facilitating the testing of MLLM. For instance, in the context of a tumor type classification task, <Task>represents “Tumor Type”, while <label 1>, <label 2>, and <label 3>are “Invasive ductal carcinoma”, “Invasive lobular carcinoma”, and “Other Type”, respectively, enabling the generation of relevant QA pairs.
Table 9: Prompt for Converting Labels into QA Pairs
[Label Transformation Prompt]
6.4 Multimodal Dataset Comparsion
Recently, several multimodal pathology datasets have been introduced for pathology applications. However, these datasets are often constrained in both scope and scale, as they primarily focus on either patch-level analysis or limited available data. In contrast, our proposed SlideInstruction and SlideBench, provided as open-source resources, significantly expand the dataset size while enhancing its versatility, as shown in Tab. 10.
@@ 6.4 Multimodal Dataset Comparsion
Dataset Level Data Type Curation Type Scope Number Availability
PathChat [21] Patch Patch and Q/A pairs Human+GPT - 257,004
Quilt-Instruct [23] Patch Patch and Q/A pairs GPT - 107,131
WSI-VQA [8] Slide WSI and Q/A pairs GPT - 8,672
PathText [7] Slide WSI-Caption pairs GPT - 9,009
HistGen [13] Slide WSI-Reports pairs GPT - 7,753
Prov-Path [31] Slide WSI-Reports pairs GPT - 17,383
CR-PathNarratives [34] Slide WSIs with annotations Human - 174
PathAlign [3] Slide WSI-Reports pairs Human - 354,089
@@
7 Experiment
7.1 Computational Cost Analysis
7.2 Implementation Details
We preprocessed each WSI by segmenting it into 224 × 224 nonoverlapping patches at a 20× magnification level, excluding background regions. We implemented our model using the Xtuner [11] toolkit and trained it across two stages on 8 × NVIDIA A100 GPUs. The training process consists of an alignment phase followed by instruction fine-tuning: Stage 1: We freeze the LLM and train the Projection and Slide Encoder with WSI-caption data for 3 epochs, using a learning rate of 0.001. Stage 2: We unfreeze the LLM, Slide Encoder, and Projection, training the model on WSI instruction-following data for 1 epoch, with a learning rate of 0.00002. Both stages are optimized using AdamW.
7.3 Ability Showcase
7.3.1 Captioning Ability
The examples shown in  Fig. 7 illustrate the capability of our model, SlideChat, to effectively perform whole-slide image captioning tasks. SlideChat demonstrates its proficiency in generating detailed and contextually accurate summaries for complex pathological whole-slide images, accurately capturing key clinical findings and pathological features. Whether summarizing broad findings, explaining pivotal details, or highlighting core results, SlideChat showcases an advanced understanding of whole-slide images, providing concise yet informative reports that align with clinical terminology and expectations.
7.3.2 VQA Ability
Fig. 8 showcases the conversational examples of SlideChat, demonstrating its ability to accurately answer a range of questions based on WSIs, covering diverse aspects such as histological classifications, tumor grading, lymph node involvement, and treatment decisions. SlideChat effectively interprets complex pathological data, engages in nuanced question-and-answer exchanges, and delivers clinically relevant responses. This reflects its potential as an intelligent assistant capable of supporting pathologists in diagnostic decision-making by providing insightful, context-aware dialogue grounded in visual pathology data.
7.3.3 Comparing Model Outputs
Fig. 9 presents a comparative analysis of the outputs from SlideChat and other models within SlideBench. The examples illustrate SlideChat’s remarkable capacity to precisely classify tumors, identify distinct histological features, and describe the structural organization of tumor cells from WSIs. SlideChat demonstrates a unique proficiency in capturing both local and global features—seamlessly integrating detailed microscopic characteristics with broader contextual understanding to deliver accurate and clinically meaningful interpretations. In contrast, existing models are limited to processing small pathology images, often yielding ambiguous or incorrect classifications. This underscores SlideChat’s advanced capability in comprehending whole-slide images by incorporating both intricate details and a comprehensive visual perspective.
7.4 Detailed Test Performance
@@
LLaVA-Med 50.04 40.63 40.38 56.95 47.34
Quilt-LLaVA 65.26 54.04 50.66 59.55 57.76
GPT-4o Slide (T) 37.07 38.76 39.93 37.60 38.28
@@
LLaVA-Med 37.25 28.57 30.41 20.71 47.27 32.78
Quilt-LLaVA 40.74 39.3 32.32 28.96 39.52 35.96
GPT-4o Slide (T) 22.95 26.76 18.06 21.06 27.82 23.10
@@
LLaVA-Med 62.04 53.98 53.04 26.54 47.96
Quilt-LLaVA 64.79 40.42 63.40 43.06 53.07
GPT-4o Slide (T) 50.00 50.08 44.16 32.64 43.42
@@
LLaVA-Med 23.95 36.62 40.19 50.76 23.72 18.99 15.05 30.10
Quilt-LLaVA 77.14 68.58 42.63 58.17 23.18 18.23 19.82 44.43
GPT-4o Slide (T) 0 0 0 0 0 0 0 0
@@
7.4.1 Performance on SlideBench-VQA (TCGA)
@@
7.4.2 Performance on SlideBench-VQA (BCNB)